OpenStudy (anonymous):

im not like this but I need help I have to finish my class&I just got called into work so I need to go my test and exam really quick I need your help for the answers please!!! A banker’s compensation consists of a weekly salary as well as a commission for each account opened. The banker earns commission at a rate of $10/account. In one particular week, the banker opened 15 accounts and brought home a total pay of $830. Which of the following equations describes the pay for the banker in terms of accounts opened? A. y = 10x + 15 B. y = 10x + 68 C. y = 15x + 680 D. y= 15x + 830

OpenStudy (anonymous):

I believe it's D. y= 15(amount of accounts opened) plus the base pay of 830.

OpenStudy (anonymous):

Scratch that. It's C. His TOTAL salary is 830, but his base pay is 680.

OpenStudy (anonymous):

What is the equation of the line that passes through the points (4/5, 1/5) and (1/2, 3/2? A. y = –x – 1 B. y = –x + 1 C. y=-13/3 x + 11/3 D. y= 7/3 x +1/3

OpenStudy (anonymous):

@Aydia thankyou!

OpenStudy (anonymous):

Here's for the second question. y=-13/3x+11/3

OpenStudy (anonymous):

What is the slope of a line that is perpendicular to 6x + 3y = –9? A.2 B.1/2 C.-2 D.-1/2

OpenStudy (anonymous):

The next one is B.

OpenStudy (radar):

Are you sure it is B? Whats wrong with D.

OpenStudy (anonymous):

I'm fairly sure. D is negative when it needs to be positive.

OpenStudy (radar):

I think A is the answer.

OpenStudy (anonymous):

What is the equation of the line that has a slope of 4/5 and a y-intercept of –1? A.y=4/5x-1 B.y=4/5x+1 C. y=-x+4/5 D.y=-x-4/5

OpenStudy (anonymous):

thankyou both you don't know how much of a life saver you are!! seriously

OpenStudy (anonymous):

I don't like numbers 😒

OpenStudy (radar):

A perpendicular line would have the negative reciprocal of the slope of the line which it is perpendicular to. Isn't that slope a - 1/2, wouldn't the negative reciprocal be a +2?

OpenStudy (radar):

Never mind I am wrong

OpenStudy (anonymous):

What is the equation of the line that passes through (4, 2) and is parallel to 3x – 2y = –6? A. y=-3/2 x+8 B.y=3/2x-4 C.y=1/2x D.y=2/3x-4

OpenStudy (radar):

the slope of the given line 6x + 3y = -9 is -2 the negative reciprocal is as you say a 1/2

OpenStudy (anonymous):

thankyou!!

OpenStudy (anonymous):

The equation of a line is y=1/4x-2. What is the equation of the line that is parallel to the first line and passes through (4, –2)? A.y=-4x+14 B.y=-1/4x-1 C.y=-4x+3 D.y=1/4x-3

OpenStudy (anonymous):

Two lines are parallel. The equation of the first line is y = –3x + 5. Which of the following cannot be the equation of the second line? A. y = –3x + 2 B. x – 3y = –3 C. 3x + y = –6 D. 6x + 2y = 2